OpenStudy (anonymous):

I would like help to find a polynomial. the question says the function F(x) = tan(x) where x in radians is to be approximated in the interval [0.2,1] to the polynomial p(x) = ax +bx^3 +cx^5 write down three linear equations when the polynomial is exact at the points x = 0.2, 0.7, 1

OpenStudy (anonymous):

Write three equations as \[ \tan (u) = a u + b u^3 + c u^5\\ \text { for } u=.2, .7, 1 \]

OpenStudy (anonymous):

\[ 0.20271=0.2 a+0.008 b+0.00032 c \\ 0.842288=0.7 a+0.343 b+0.16807 c\\ 1.55741= a+ b +c \]

OpenStudy (anonymous):

If you solve them, you get \[ a = 1.00226,\\ b = 0.270994,\\ c = 0.284158 \]

OpenStudy (anonymous):

thank you for this answer, where I was going wrong was for some reason I was ignoring tan(x). I couldn't see the woods for the tree,s jk67

OpenStudy (anonymous):

yw